Session: Oct 2007 (session 3)
Wavelength: 0.7 cm


To give feedback on this experiment, click here

Experiment Feedback

Noto (P. Cassaro at Wed Oct 24 10:29:13 2007) Minor failures : TSYS
The receiver works with the only left polarization, due to a problem arisen just before the session. Hence we have no right pol in the data.

Onsala (Michael Lindqvist at Thu Oct 25 09:46:37 2007) Minor failures : ANT
20m antenna *might* have been incorrectly initialized at the start of the experiment. Incorrect initialization could mean incorrect tracking. Took some time before we noticed. Due to this, the schedule was restarted around 01:00 UTC. However, later, after some testing and dicussions we have come to the conclusion that the data is *probably* OK before the restart.

Effelsberg (Bach/Polatidis at Mon Nov 5 10:52:10 2007) Minor failures : RECE
Unfortunate choice of LO settings caused probelms with phase lock. Problem was solved at about 18 UT. BBC3+4 both LCP and RCP were low.